OpenStudy (anonymous):

What is the approximate circumference of a semicircle with a radius of 35 inches? A. 1923.25 inches B. 219.8 inches C. 109.9 inches D. 54.95 inches

OpenStudy (anonymous):

circumference of a circle with radius r is \[2\pi r\] so, for a semicircle it will be \[\pi r\] For approx calculations assume \[\pi\] = 3

OpenStudy (linn99123):

so it be C right ^^?

OpenStudy (anonymous):

thank you sooo much linn